what is
how i can graph this
y=6
2x-2y=-4

Answer by mananth(16946)

y=6 is a line parallel to the x axis intercepting the y axis at 6
2x-2y=-4
/2
x-y=-2
Generate values for y from values of x
Assume any three or more values for x and plug in the equation.

x 0 -3 3 6
y= 2 -1 5 8
(x,y) (0,2) (-3,-1) (3,5) (6,8)
plot the points and draw the line


m.ananth@hotmail.ca